# What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?

### v3

@@ 3.3 What About the Timesteps and Text Prompts?
Table 3: Quantitative comparisons among different timesteps and text prompts of diffusion models on depth estimation.
@@ 3.4 How to Leverage the U-Net’s Prior Knowledge?
The significance of the denoiser cannot be overstated. However, the strategies for its utilization are worth careful consideration. Should we freeze the denoiser, utilize its intermediate features, and merely fine-tune the decoder for specific tasks? Alternatively, could we employ LoRA (Hu et al., 2022) instead of extensively fine-tuning the entire denoiser? Unfortunately, the evidence suggests that neither approach is ideal. As illustrated in table 4, freezing the denoiser significantly compromises performance. Although incorporating LoRA offers some advantages, it can not fully leverage the potential of denoiser, especially with regular LoRA ranks of 4 and 16. This limitation likely stems from the substantial differences between the noise-to-image denoising process and the image-to-perception prediction task.
@@
Fine-tuning the denoiser is preferable compared to either merely utilizing its intermediate features or training it with LoRA.
@@ 3.6 Summary of the obervations
Based on the preceding analysis, an effective approach to leveraging the prior knowledge of diffusion models is to use them as single-step deterministic perception estimators. This can be done with either a VAE decoder or a customized lightweight decoder. Additionally, employing pixel-specific customized loss functions can further enhance detail and overall performance. We compare our deterministic single-step perception method with previous multi-step paradigms in fig. 1. In the following section, we extend these findings to a broader set of visual perception tasks, including surface normal estimation, semantic image segmentation, dichotomous image segmentation, and image matting.
@@ 4.1 Geometric Estimation
Table 6: Quantitative comparison of affine-invariant depth estimation on five zero-shot datasets.
@@
Table 7: Quantitative comparison of surface normal estimation on three zero-shot datasets. We evaluate mean $\downarrow$, median error $\downarrow$ (med.), and the percentages of pixels $\uparrow$ with five thresholds.
@@
Quantitative Evaluation. Quantitative results on monocular depth estimation and surface normal estimation are shown in table 6 and table 7, respectively. Even trained on limited synthetic datasets only, our GenPercept shows much robustness and achieves promising performance on diverse unseen scenes. For monocular depth models, we train them with pixel-wise MSE loss, scale-shift-invariant loss (Ranftl et al., 2020), and gradient loss (Ranftl et al., 2020). Furthermore, our disparity model (inverse of the depth) shows much better performance on datasets with outdoor scenes, such as KITTI and DIODE, but less performance on indoor datasets. By replacing the VAE decoder with a lightweight DPT head, GenPercept can infer faster without bearing the performance penalty. For surface normal estimation, the image angular loss brings significant performance improvement thanks to our one-step estimation paradigm.
Qualitative Results. Qualitative visualizations are shown in fig. 3. We observe excellent generalization of our models in that they can estimate accurate geometric information and promising details not only on diverse real scenes and synthetic scenes, but also on comics, color drafts, and even some sketches.
@@ 4.2 Image Segmentation
Table 9: Quantitative results of semantic segmentation performance on Hypersim and ADE20k datasets.
@@
(Hypersim) (ADE20K)
Our GenPercept 46.0 34.1
Our GenPercept 52.9 38.3
Table 10: Quantitative comparisons of image matting on the P3M-500-NP test set.
Method SAD $\downarrow$ MAD $\downarrow$ MSE $\downarrow$ CONN $\downarrow$
HATT (Qiao et al., 2020) 30.35 0.0176 0.0072 27.42
@@
Quantitative results of dichotomous image segmentation are shown in table 8, respectively. We only show partial results due to paper page limitations, full comparisons are accessible in the supplementary material. Our proposed model outperforms all other methods by a significant margin on this challenging dataset across most evaluation metrics. As shown in Fig. 5, the segmentation results of our approach are more refined, providing a more detailed foreground mask. For thin lines and meticulous objects that are difficult for previous methods to process, our method can also output accurate segmentation results.
@@
Another choice involves using a customized segmentation head. We incorporate a custom segmentation head, namely UperNet (Xiao et al., 2018b), onto the multi-level features extracted by UNet, and the quantitative results are presented in table 10. In addition to evaluating on Hypersim (Roberts et al., 2021), we also test the model’s zero-shot performance on a subset of the ADE20k (Zhou et al., 2017) validation set, which contains overlapping classes. The results demonstrate the model’s capacity for zero-shot generalization.
@@ 4.3 Image Matting
Task Definition. Image matting aims to extract the foreground, background, and alpha mask from an image. Traditional approaches depend on supplementary inputs that delineate foreground, background, and ambiguous areas to reduce uncertainty. Automatic image matting seeks to remove this dependency by directly estimating these components from the image alone.
Quantitative and Qualitative Results. We evaluate metrics including the sum of absolute differences (SAD), mean squared error (MSE), mean absolute difference (MAD), gradient (Grad.), and Connectivity (Conn.) on the P3M-500-NP test set. SAD and MAD measure the mean L1 distance between predictions and ground truth labels. MSE and CONN focus on L2 distance and connectivity that better reflects human intuition. As shown in table 10, our GenPercept is comparable with the state-of-the-art methods. Quantitative results of image matting are shown in fig. 6.
@@ 5 Related work
Vision Pre-Training. Models pretrained on large-scale datasets possess powerful feature extraction capabilities, enabling them to be effectively transferred to a wide range of visual tasks. For instance, the ResNet (He et al., 2016) model pretrained on ImageNet (Russakovsky et al., 2015) can be fine-tuned and applied to perception tasks. By means of contrastive learning, MoCo (He et al., 2020) and CLIP (Radford et al., 2021) acquire rich visual and semantic representations, leveraging their advantages in joint visual and semantic modeling to enhance the performance of multimodal tasks. DINO (Zhang et al., 2022b), through self-distillation, endows Vision Transformer and convolutional networks with comparable visual representation quality and demonstrates that self-supervised ViT representations contain explicit semantic segmentation information. In our work, we leverage Stable Diffusion (Rombach et al., 2022) as a prior for scene understanding and transfer it to various perception tasks.
@@ 6 Conclusion
In this work, we introduce GenPercept, an embarrassingly straightforward yet powerful approach to re-use the off-the-shelf UNet trained using diffusion processes. In essence, we initialize our image perception networks using pre-trained diffusion models trained on the extensive LAION dataset, without making many modifications. GenPercept demonstrates the capability to effectively leverage pre-trained diffusion models across a range of downstream dense perception tasks. We contend that our proposed methodology provides an efficient and potent paradigm for harnessing the capabilities of pre-trained diffusion models in dense visual perception tasks.
Supplementary Materials of
What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?
@@ A.1 Stochastic Multi-Step Generation
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$, and $\beta_{s}$ is sampled from a variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$. Then, the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the velocity of “v-prediction” Salimans & Ho (2021) from a timestep $t$ and the corresponding input $\mathbf{z}_{t}$. During training, the parameters of VAE are frozen, and only the denoiser $\bm{v}_{\theta}$ is fine-tuned. The timestep is uniformly sampled from 1 to $T$.
@@
where the denoising process first computes the estimated noise $\hat{\bm{\bm{\epsilon}}}_{t}$ and the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}$ of timestep $t$ from the current latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{t}$ and the predicted velocity $\bm{v}_{\theta}(\mathbf{z}_{t},t)$. Then, it adds the computed noise $\hat{\bm{\bm{\epsilon}}}_{t}$ back to $\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}$ to get the latent code of timestep $t-1$. After repeating it for $T$ times, the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}$ is computed and sent to the VAE decoder $\mathcal{D}$ and estimate the target label $\hat{\mathbf{y}}$. During inference, $m$ randomly sampled noises are introduced to estimate $m$ different predictions, and they are averaged with an ensemble process to reduce the randomness of prediction and improve performance.
@@ A.2 Deterministic Multi-Step Generation
Figure 7: Qualitative visualization of ablation study.
@@ A.3 GenPercept: Deterministic One-step Perception
In the main paper, we suppose that blending with the RGB image is much simpler compared to Gaussian noise, and we increase the blending proportion of the RGB image to make it more difficult. This is realized by increasing the initial value $\beta_{0}$ and final values $\beta_{T}$ of the DDPM schedule. When all the beta values $\{\beta_{t}\}_{t=1}^{T}$ are increased to 1, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}=0$, the formulation of eq. 4 to eq. 6 can be greatly simplified as follows.
@@
Appendix B More Visualization Analysis

### v4

@@ 3.3 What About the Timesteps and Text Prompts?
Table 3: Quantitative comparisons among different timesteps and text prompts on depth estimation.
@@ 3.4 How to Leverage the U-Net’s Prior Knowledge?
The significance of the denoiser cannot be overstated. However, the strategies for its utilization are worth a careful study. Should we freeze the denoiser, utilize its intermediate features, and merely fine-tune the decoder for specific tasks? Alternatively, can we employ LoRA (Hu et al., 2022) instead of extensively fine-tuning the entire denoiser? Unfortunately, the evidence suggests that neither approach is ideal. As illustrated in Table 4, freezing the denoiser significantly compromises performance. Although incorporating LoRA offers some advantages, it may not fully leverage the potential of denoiser, especially with regular LoRA ranks of 4 and 16. This limitation likely stems from the substantial differences between the noise-to-image denoising process and the image-to-perception prediction task.
@@
Fine-tuning the denoiser appears to be preferable for achieving better results, compared to either merely utilizing its intermediate features or training a LoRA.
@@ 3.6 Summary of the obervations
Based on the preceding analysis, an effective approach to leveraging the prior knowledge of diffusion models is to use them as single-step deterministic perception estimators. This can be done with either a VAE decoder or a customized lightweight decoder. Additionally, employing pixel-specific customized losses can further enhance detail and overall performance. We compare our deterministic single-step perception method with previous multi-step paradigms in Fig. 1. In the following section, we extend these findings to a broader set of visual perception tasks, including surface normal estimation, semantic image segmentation, dichotomous image segmentation, and image matting.
@@ 4.1 Geometric Estimation
For geometry evaluation, the ensemble size, inference resolution, valid evaluation depth range (specific for depth estimation), and evaluation average paradigm (average by pixels or average by the number of images) can be different for each method. To compare these approaches fairly, we follow the open-source evaluation code of Marigold (Ke et al., 2024) for depth and DSINE (Bae & Davison, 2024a) for surface normal, and evaluate the performance of partial existing SOTA methods with their officially released model weights. They are labeled with in the Table.
@@
Table 6: Quantitative comparison of affine-invariant depth estimation on five zero-shot datasets. Part of the reported results () are evaluated following the evaluation protocol of Marigold by ourselves.
@@
Table 7: Quantitative comparison of surface normal estimation on three zero-shot datasets. We evaluate mean $\downarrow$, median error $\downarrow$ (med.), and the percentages of pixels $\uparrow$ with five thresholds. Part of the reported results () are evaluated following the evaluation protocol of DSINE by ourselves.
@@
Quantitative Evaluation. Quantitative results on monocular depth estimation and surface normal estimation are shown in Table 6 and Table 7, respectively. Even trained on limited synthetic datasets only, our GenPercept shows much robustness and achieves promising performance on diverse unseen scenes. For monocular depth models, we train them with pixel-wise MSE loss, scale-shift-invariant loss (Ranftl et al., 2020), and gradient loss (Ranftl et al., 2020). Furthermore, our disparity model (inverse of the depth) shows much better performance on datasets with outdoor scenes, such as KITTI and DIODE, but less performance on indoor datasets. Therefore, we suggest adopting the depth model for indoor scenes and the disparity model for outdoor scenes experimentally. By replacing the VAE decoder with a lightweight DPT head (Ranftl et al., 2021), GenPercept can infer faster without bearing the performance penalty. For surface normal estimation, the image angular loss brings significant performance improvement thanks to our one-step estimation paradigm.
Qualitative Results. Qualitative visualizations are shown in Fig. 3. We observe excellent generalization of our models in that they can estimate accurate geometric information and promising details not only on diverse real and synthetic scenes, but also on comics, color drafts, and even sketches.
@@ 4.2 Image Segmentation
Our Genpercept (infer. at 1024px) 0.877 0.859 0.035 0.887 0.941 1262 0.874 0.858 0.041 0.874 0.947 3321 0.875 0.856 0.036 0.885 0.939 1176
Table 9: Quantitative results of semantic segmentation on Hypersim and ADE20k.
@@
Dataset (Hypersim) (ADE20K)
GenPercept (Train UperNet) Hypersim 46.0 34.1
GenPercept (Train U-Net + UperNet) 52.9 38.3
Mask2Former R50 ADE20K - 47.2
Mask2Former Swin-T - 47.7
Mask2Former Swin-L - 56.4
GenPercept (Train U-Net + UperNet) - 50.2
Table 10: Quantitative comparisons of image matting on the P3M-500-NP and AIM500.
Method Test Dataset SAD $\downarrow$ MAD $\downarrow$ MSE $\downarrow$ CONN $\downarrow$
HATT (Qiao et al., 2020) P3M-500-NP 30.35 0.0176 0.0072 27.42
@@
Quantitative results of dichotomous image segmentation are shown in Table 8. We only show partial results due to paper page limitations, full comparisons are accessible in the supplementary material. GenPercept outperforms methods like HySM (Nirkin et al., 2021) and IS-Net (Qin et al., 2022) on this challenging dataset across most evaluation metrics, but there exists room for further improvement compared to SoTA methods like MVANet (Yu et al., 2024). As shown in Fig. 5, our approach provides a detailed foreground mask. For thin lines and meticulous objects that are difficult for previous methods to process, our method can also output accurate segmentation results.
@@
Another choice involves using a customized segmentation head. We incorporate a custom segmentation head, namely UperNet (Xiao et al., 2018b), onto the multi-level features extracted by UNet. For the UperNet segmentation head, we follow the traditional semantic segmentation format to use n-channel output, where n is the number of categories. The quantitative results are presented in Table 10, we test the model’s performance on Hypersim (Roberts et al., 2021) and zero-shot ability on a subset of the ADE20k (Zhou et al., 2017) validation set, which contains overlapping classes. Besides, we compare with Mask2Former (Cheng et al., 2022) by training on ADE20K. GenPercept outperforms ResNet50 (He et al., 2016) and Swin-T (Liu et al., 2021) of Mask2Former but achieves lower performance than Swin-L (Liu et al., 2021).
@@ 4.3 Image Matting
Task Definition. Image matting aims to extract the foreground, background, and alpha mask from an image. Traditional approaches depend on supplementary inputs that delineate foreground, background, and ambiguous areas to reduce uncertainty. Automatic image matting seeks to remove this dependency by directly estimating these components from the image alone. The implementation details can be found in the supplementary material.
Quantitative and Qualitative Results. We evaluate metrics including the sum of absolute differences (SAD), mean squared error (MSE), mean absolute difference (MAD), gradient (Grad.), and Connectivity (Conn.) on the P3M-500-NP test set. SAD and MAD measure the mean L1 distance between predictions and ground truth labels. MSE and CONN focus on L2 distance and connectivity that better reflects human intuition. As shown in Table 10, our GenPercept is less accurate compared with the state-of-the-art methods. However, when transferring the human image matting ability to general image matting tasks, GenPercept achieves much better performance. It proves the robustness brought by the prior knowledge of diffusion models pre-trained on the LAION dataset. Quantitative results of image matting are shown in Fig. 6. Please see supplementary for more visualization.
@@ 5 Related work
Vision Pre-Training. Models pretrained on large-scale datasets possess powerful feature extraction capabilities, enabling them to be effectively transferred to a wide range of visual tasks. For instance, the ResNet (He et al., 2016) model pretrained on ImageNet (Russakovsky et al., 2015) can be fine-tuned and applied to perception tasks. By means of contrastive learning, MoCo (He et al., 2020) and CLIP (Radford et al., 2021) acquire rich visual and semantic representations, leveraging their advantages in joint visual and semantic modeling to enhance the performance of multimodal tasks. DINO (Caron et al., 2021), through self-distillation, endows Vision Transformer and convolutional networks with comparable visual representation quality and demonstrates that self-supervised ViT representations contain explicit semantic segmentation information. DINOv2 (Oquab et al., 2024) leverages self-supervised learning on a large curated dataset and exhibits remarkable zero-shot generalization capabilities across computer vision tasks at both image and pixel levels, including classification, semantic segmentation, and depth estimation. In our work, we leverage Stable Diffusion (Rombach et al., 2022) as a prior for scene understanding and transfer it to various perception tasks.
@@ 6 Conclusion
In this work, we introduce GenPercept, an embarrassingly straightforward yet powerful approach to re-use the off-the-shelf UNet trained using diffusion processes. GenPercept demonstrates the capability to effectively leverage pre-trained diffusion models across a range of downstream dense perception tasks. We contend that our proposed methodology provides an efficient and potent paradigm for harnessing the capabilities of pre-trained diffusion models in dense visual perception tasks.
For future research, we suggest investigating the impact of scaling up the volume of fine-tuning data and exploring the key components of pre-training by applying alternative self-supervised pre-training methods on the LAION dataset, such as Masked Autoencoders (MAE) or Contrastive Language-Image Pretraining (CLIP). It will be helpful to clarify whether the highly detailed visual predictions produced by existing diffusion models are primarily driven by the extensive LAION dataset or the diffusion pretraining paradigm itself.
Acknowledgments
Y. Ge is with The University of Adelaide and his contribution was made when visiting Zhejiang University. C. Shen is the corresponding author.
@@ A.1 Stochastic Multi-Step Generation
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$, and $\beta_{s}$ is sampled from a variance scheduler $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$. The scheduler is parameters with two hyperparameters $\beta_{start}$ and $\beta_{end}$, which defines the $\beta_{t}$ values of t=0 and t=1000, respectively. For a casual timestep $s$, $\beta_{s}$ is computed by linearly interpolating between $\sqrt{\beta_{start}}$ and $\sqrt{\beta_{end}}$, then squaring each interpolated value. The formulation is as follows.
$\begin{gathered}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta_{s}=(\sqrt{\beta_{start}}+\frac{s}{T}(\sqrt{\beta_{end}}-\sqrt{\beta_{start}}))^{2},~{}~{}s=[1,...,T]}\end{gathered}$ (5)
Then, the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the “v-prediction” (Salimans & Ho, 2021) from a timestep $t$ and the corresponding input $\mathbf{z}_{t}$. During training, the parameters of VAE are frozen, and only the denoiser $\bm{v}_{\theta}$ is fine-tuned. The timestep is uniformly sampled from 1 to $T$.
@@
where the denoising process first computes the estimated noise $\hat{\bm{\bm{\epsilon}}}_{t}$ and the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}$ of timestep $t$ from the current latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{t}$ and the predicted velocity $\bm{v}_{\theta}(\mathbf{z}_{t},t)$. Then, it adds the computed noise $\hat{\bm{\bm{\epsilon}}}_{t}$ back to $\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}$ to get the latent code of timestep $t-1$. After repeating it for $T$ times, the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}$ is computed and sent to the VAE decoder $\mathcal{D}$ and estimate the target label $\hat{\mathbf{y}}$. During inference, $m$ randomly sampled noises are introduced to estimate $m$ different predictions, and they are averaged with an ensemble process (Ke et al., 2024) to reduce the randomness of prediction and improve performance.
@@ A.2 Deterministic Multi-Step Generation
Table 11: Runtime comparison of three diffusion for perception pipelines on an RTX 4090 GPU.
Experimental Setting Ensemble Denoise Steps Inference Time GPU Memory
Stochastic Multi-step Generation (w. ensemble) 10 10 $\sim$5.74s 16GB
Stochastic Multi-step Generation (w/o ensemble) 1 10 $\sim$0.79s 6.95GB
Deterministic Multi-step Generation 1 10 $\sim$0.79s 6.95GB
Deterministic One-step Inference (Ours) 1 1 $\sim$0.34s 6.95GB
Deterministic One-step Inference + DPT head (Ours) 1 1 $\sim$0.24s 6.32GB
Metric3Dv2 (Hu et al., 2024) 1 1 $\sim$0.25s 2.63GB
DepthAnythingv2 (Yang et al., 2024b) 1 1 $\sim$0.07s 2.82GB
DSINE (Bae & Davison, 2024a) 1 1 $\sim$0.18s 2.23GB
Marigold (Ke et al., 2024) 1 10 $\sim$0.79s 6.95GB
GeoWizard (Fu et al., 2024b) 1 1 $\sim$1.32s 6.81GB
DepthFM (Gui et al., 2024) 1 2 $\sim$0.41s 6.97GB
Our GenPercept (DPT head) 1 1 $\sim$0.24s 6.32GB
Table 12: The impact of training data volume on affine-invariant monocular depth estimation.
Amount of Data KITTI NYU ScanNet DIODE ETH3D
[…]
90K (1/1) 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956
45K (1/2) 0.101 0.902 0.056 0.964 0.058 0.963 0.311 0.764 0.070 0.955
22.5K (1/4) 0.109 0.884 0.056 0.963 0.059 0.963 0.322 0.754 0.073 0.950
11.2K (1/8) 0.117 0.866 0.060 0.962 0.065 0.957 0.324 0.753 0.076 0.943
5.6K (1/16) 0.117 0.868 0.063 0.958 0.068 0.952 0.331 0.744 0.084 0.932
@@ A.3 GenPercept: Deterministic One-step Perception
In the main paper, We set the ($\beta_{start}$, $\beta_{end}$) values to 1, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}=0$, the formulation of Eq. 8 to Eq. 10 can be greatly simplified as follows.
@@
Appendix B A Difference Between Image Generation and Visual Perception
In text-guided image generation, a single textual input can correspond to an immense variety of potential images. This inherent uncertainty makes generating a high-quality image directly from random noise in a single step extremely challenging. Therefore, the multi-step generation enables the model to incrementally remove noise, progressively refining details and textures at each stage, which effectively simplifies the task. However, visual perception tasks conditioned on an RGB image are deterministic without any randomness, and such an easy injective mapping can be estimated with a one-step inference process, as most of the traditional visual perception methods do.
While Marigold series algorithms aim to leverage diffusion models’ ability of generating highly detailed images to enhance visual perception with precise details, reformulating straightforward deterministic tasks as a denoising process can further simplify this problem, enforcing the network to exploit ”shortcuts”, as described in Section 3.1 of the main paper and illustrated in Fig. 8.
Appendix C Runtime Analysis
In this section, we quantitatively analyze the inference times of the three aforementioned pipelines, as summarized in Table 11. The runtime is evaluated by averaging the inference times over 100 images with a resolution of $768\times 768$, using an RTX 4090 GPU. For ”Stochastic Multi-step Generation” methods (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024), such as Marigold (Ke et al., 2024), they rely on an ensemble process where multiple inferences are performed with varying random noise inputs to mitigate uncertainties introduced by Gaussian noise. Consequently, this approach is computationally expensive. On the other hand, ”Deterministic Multi-step Generation” methods (Lee et al., 2024) involve multiple denoising steps, which significantly reduce inference efficiency.
In contrast, our proposed one-step inference paradigm demonstrates a runtime that is 94% and 57% less than those of multi-step methods with ensemble and without ensemble, respectively. Furthermore, by incorporating a customized head, such as the DPT head (Ranftl et al., 2021), both runtime and GPU memory requirements are further reduced by 27% without compromising performance, maintaining a competitive level of accuracy and robustness.
Compared to existing state-of-the-art diffusion-based methods, our proposed GenPercept achieves a notable improvement in inference speed, attributed to the innovative one-step inference paradigm and the customized head. While our method demonstrates inference speeds comparable to Metric3Dv2 (Hu et al., 2024) and DSINE (Bae & Davison, 2024a), it falls behind DepthAnythingV2 (Yang et al., 2024b). Note that the superior performance of DepthAnythingV2 is facilitated by its training on a relatively lightweight model, bolstered by extensive labeled and unlabeled datasets, and supported by substantial computational resources distributed across multiple GPUs.
Figure 7: Qualitative visualization of ablation study.
Appendix D The Impact of Data Volume
With only around 50K Hypersim (Roberts et al., 2021) and 40K Virtual KITTI (Cabon et al., 2020) fine-tuning data, GenPercept can generalize well to diverse tasks and datasets. How much data is needed for transferring at least? We gradually reduce the amount of training data. As shown in Table 12, less training data results in slightly worse performance, but GenPercept still shows much robustness to the data volume.
Appendix E More Visualization Analysis